Betty invests $12,000 in two plans. One pays 6% interest and the other pays 9% interest. If the total interest is $828, how much was invested in each plan?
Answer by ewatrrr(24785)

Hi,
Let x represent the amount invested at 6%. then (12,000-x) would be the amount remaining to be invested at 9%
.
question states;
.06x + .09(12,000-x) = $828
.
simplify
.06s + 1080 - .09x = $828
1080-828 = .03x
252 = .03x
$8400 = x
.
$8400 at 6% and $3600 at 9%